Show that the quadrilateral with vertices (0,1)(4,2)(3,6)(-5,4) has two right angles. Find the area
Answer by richwmiller(17219)

(0,1)(4,2)(3,6)(-5,4)
(0,1)(4,2)(3,6) forms right angle
(4,2)(3,6) y = 18-4 x
(0,1)(4,2) y = x/4+1
slopes are -4 and 1/4
-4*1/4=-1 they are perpendicular and form a right angle
(4,2)(3,6)(-5,4) forms right angle
(3,6)(-5,4) y = x/4+21/4
(4,2)(3,6) y = 18-4 x
slopes are -4 and 1/4
-4*1/4=-1 they are perpendicular and form a right angle
area 25.5